A solution contains 0.440 M potassium nitrite and 0.102 M nitrous acid.

The pH of this solution is .
Answer #1

mixture of HNO2 and KNO2 act as acidic buffer

pH = pKa + log [NaNO2] / [HNO2]

pKa = HNO2 = 3.15

pH = 3.15 + log [0.440]/[0.102]

pH = 3.15 + 0.63

pH = 3.78